# GuidedSampling: Steering LLMs Towards Diverse Candidate Solutions at Inference-Time

### v1

@@ 2 Related Works
Chain-of-thought (CoT) and its variants (Wei et al., 2022; Kojima et al., 2022) showed that guiding LLMs to produce intermediate reasoning steps during inference boosts performance on complex tasks such as mathematical and commonsense reasoning. However, as reasoning chains become longer, CoT suffers from error propagation due to complex calculations (Chen et al., 2022). To mitigate this, new methods have been proposed, e.g., Self-Consistency, which samples multiple CoT from LLM and then selects the most consistent final answer through majority voting (Wang et al., 2022). Building upon these ideas, better search algorithms, such as the tree-of-thought (ToT) (Yao et al., 2023), MCTS (Zhang et al., 2024), and REBASE (Wu et al., 2024), have been proposed, which enable LLMs to perform more deliberate problem-solving by exploring multiple reasoning paths in a tree structure. Several agentic systems (Parmar et al., 2025; Estornell & Liu, 2024) have shown that performing multi-agent debate at inference before generating a final solution improves performance. Furthermore, recent work (Muennighoff et al., 2025) has extended the ‘thinking’ of models by introducing special tokens such as “wait” to improve performance. Finally, Ghosal et al. (2025) has shown that simply sampling from a model repeatedly outperforms such approaches. In contrast to prior methods, GuidedSampling generates a diverse set of samples with lower inference-time cost than tree search (Yao et al., 2023), while achieving greater diversity than standard sampling approaches. Parallel to our work, Wang et al. (2025) proposed RandIdeaInjection, which first generates a list of distinct ideas and then injects the generated list into the generation process to produce the final response. GuidedSampling, on the other hand, works in an iterative loop of generating concepts, adding them individually to generate the final output.
@@
Recent works have explored leveraging advanced inference strategies for both generating high-quality synthetic training data and for fine-tuning models to improve their performance. For instance, Self-Taught Reasoner (STaR) (Zelikman et al., 2022) is an iterative method where an LLM is prompted to generate CoT rationales; those rationales that lead to correct answers are then used as high-quality synthetic data to fine-tune the model, while those which lead to incorrect answers are passed back to model for refinement along with the correct final answer, effectively bootstrapping its reasoning abilities from a small initial set. Similarly, ReSTEM (Singh et al., 2023), building on principles of reinforced self-training (ReST), employs an iterative Expectation-Maximization-like framework. It uses Best-of-N (BoN) sampling to generate multiple candidate solutions for problems and then refines the model by training on this synthetically generated data. Chow et al. (2024) and Tang et al. (2025) developed reinforcement learning methods that directly optimize for pass@k metrics and majority voting performance, leading to significant gains in reasoning and code generation. Other methods, such as multi-agent fine-tuning (Subramaniam et al., 2025), train diverse agent models through debate and voting, while Gui et al. (2024) introduced BoNBoN Alignment, distilling the BoN distribution into a single model. While these strategies improve pass@k, they often do not explicitly manage the trade-off between exploration and generation. In contrast, our proposed GuidedSampling method introduces a structured exploration phase during training, explicitly balancing diversity and quality, and models fine-tuned with our trajectories outperform those trained using methods like BoN, STaR, or ToT.
@@ 3.3 Theoretical Bounds for GuidedSampling
  • $\mathcal{I}(y;c\;|\;x)$: sample-wise mutual information between $y$ and $c$ conditional on $x$. This represents the additional information contributed by the concept $c$ in predicting the answer $y$.
To elaborate on the performance bounds of GuidedSampling, we make the following assumption:
@@
For any “informative” concept $c\in\mathcal{C}_{inf}$, conditioning on it strictly increases the probability of generating a correct solution $y^{*}$. That is, there exists an amplification factor $k_{c}>1$ such that:
$\mathcal{I}(y;c\;|\;x)>0$
$\log\pi_{base}(y^{*}\;|\;x,c)\;-\;\log\pi_{base}(y^{*}\;|\;x)>0$
$\log\pi_{base}(y^{*}\;|\;x,c)>\log\pi_{base}(y^{*}\;|\;x)$
$\pi_{base}(y^{*}\;|\;x,c)=k_{c}\cdot\pi_{base}(y^{*}\;|\;x)$ (1)
Following the above assumption, we now state our main theorem:
@@
The condition derived in Theorem 1 provides a formal basis for when GuidedSampling outperforms RS. We detail the proof in Appendix A. In practice, this condition is satisfied if one or more of the following pathways hold:
@@ 3.4 Post-Training using GuidedSampling
Synthetic data has become an increasingly effective tool for enhancing the reasoning capabilities of LLMs (Gupta et al., 2023; Mitra et al., 2024; Chaudhary et al., 2023). In particular, inference-time algorithms are valuable for generating such data when the correctness of the final solution can be programmatically verified (Zelikman et al., 2022; Arora & Zanette, 2025; Shao et al., 2024). We demonstrate that GuidedSampling can serve not only as an effective inference-time strategy but also as a powerful synthetic data generation mechanism.
@@ 5 Results and Discussion
Figure 4: GuidedSampling forces exploration during inference-time, resulting in $16.01\%$ average pass@k improvement compared to repeated sampling. We observe an average improvement of $21.8\%$ on MATH, $11.87\%$ on GPQA-Diamond, $11.28\%$ on HumanEval, and $3.08\%$ on OlympiadBench. First row: For Llama-3.2-3B-Instruct, Second row: For Qwen2.5-3B-Instruct.
@@
To measure the diversity of candidate solutions, we use Qwen2.5-32B-Instruct (Yang et al., 2024) to extract the core concept or theorem. We then compute the number of distinct concepts generated. On average, RS produces 3.54, 6.72, 2.66, and 3.25 distinct concepts on MATH, GPQA-Diamond, HumanEval, and OlympiadBench, respectively. GuidedSampling improves this diversity by an average of 17.63%. We also found the diversity gains from GuidedSampling to be model-specific. We find that Llama-3.2-3B-Instruct generates $3.7\times$ more unique ideas on average compared to Qwen2.5-3B-Instruct, with this gap ranging from $2.82\times$ on GPQA‑Diamond to $5.12\times$ on HumanEval. This suggests model architecture and pretraining heavily influence the capacity for generating novel reasoning strategies. We illustrate some examples of generated concepts in Appendix C.
Figure 5: Pass@50 performance variation with different exploration (number of concepts $K$) and generation (samples per idea $M$) compute allocations, given a fixed total compute of 100 calls ($M=100/K$). Increasing exploration initially helps, but performance declines when the generation budget per idea becomes too small. At $K=0$, GuidedSampling becomes traditional RS. The first row shows results for Llama-3.2-3B-Instruct, and the second for Qwen2.5-3B-Instruct.
@@ 6 Conclusions
Appendix A Proof of Theorem 1
@@ B.2 Concept Extraction Prompt
IAA Data You are ConceptTagger, an expert that maps a worked-out solution (chain-of-thought or final answer) to the most specific mathematical or logical concept that makes the solution possible.
@@
B.3 IAA Prompt
IAA Data I have a few ideas to solve this problem.
@@ Appendix E More results using GuidedSampling
In this section, we showcase some results on additional models. As mentioned in §4, we generate 100 candidate solutions for each instance. We provide results on Phi-4-mini-instruct (Abouelenin et al., 2025) and GPT-4o-mini (Hurst et al., 2024). Due to limited resource constraints, we limit both these models to just MATH (Hendrycks et al., 2021) benchmark. Table 6 shows the pass@50 results for both the models along with the observed diversity as extracted by Qwen-3.2-32B-Instruct (Yang et al., 2024). Diversity is measured by the average number of concepts for each instance.

### v2

@@ 2 Related Works
Chain-of-thought (CoT) and its variants (Wei et al., 2022; Kojima et al., 2022) showed that guiding LLMs to produce intermediate reasoning steps during inference boosts performance on complex tasks such as mathematical reasoning. However, as reasoning chains become longer, CoT suffers from error propagation due to complex calculations (Chen et al., 2022). To mitigate this, new methods have been proposed, e.g., Self-Consistency, which samples multiple CoT and then selects the most consistent answer through majority voting (Wang et al., 2022). Building upon these ideas, better search algorithms, such as the tree-of-thought (ToT) (Yao et al., 2023), MCTS (Zhang et al., 2024), and REBASE (Wu et al., 2024), have been proposed, which enable LLMs to perform more deliberate problem-solving by exploring multiple reasoning paths in a tree structure. Several agentic systems (Parmar et al., 2025; Handa et al., 2025; Zhang et al., 2025; Kumbhar et al., 2025) have shown that performing multi-agent debate before generating a final solution improves performance. Furthermore, recent work (Muennighoff et al., 2025) has extended the ‘thinking’ of models by introducing special tokens such as “wait” to improve performance. Finally, Ghosal et al. (2025) has shown that simply sampling from a model repeatedly outperforms such approaches. In contrast to prior methods, GuidedSampling generates a diverse set of samples with lower inference-time cost than tree search (Yao et al., 2023), while achieving greater diversity than standard sampling approaches. Parallel to our work, Wang et al. (2025) proposed RandIdeaInjection, which first generates a list of distinct ideas and then injects the generated list into the generation process to produce the final response. GuidedSampling, on the other hand, works in an iterative loop of generating concepts, adding them individually to generate the final output.
@@
Recent works have explored leveraging inference strategies for generating high-quality synthetic training data and fine-tuning models to improve their performance. For instance, Self-Taught Reasoner (STaR) (Zelikman et al., 2022) is an iterative method where an LLM is prompted to generate CoT rationales; those rationales that lead to correct answers are then used as high-quality synthetic data to fine-tune the model, while those which lead to incorrect answers are passed back to model for refinement along with the correct final answer, effectively bootstrapping its reasoning abilities from a small initial set. Similarly, ReSTEM (Singh et al., 2023), building on principles of reinforced self-training (ReST), employs an iterative Expectation-Maximization-like framework. It uses Best-of-N (BoN) sampling to generate multiple candidate solutions for problems and then refines the model by training on this synthetically generated data. Some works (Chow et al., 2024; Tang et al., 2025; Rrv et al., 2025) have developed reinforcement learning methods that directly optimize for pass@k metrics, leading to significant gains in reasoning and code generation. Other methods, such as multi-agent fine-tuning (Subramaniam et al., 2025), train diverse agent models through debate and voting, while Gui et al. (2024) introduced BoNBoN Alignment, distilling the BoN distribution into a single model. While these strategies improve pass@k, they often do not explicitly manage the trade-off between exploration and generation. In contrast, our proposed GuidedSampling method introduces a structured exploration phase during training, explicitly balancing diversity and quality, and models fine-tuned with our trajectories outperform those trained using methods like BoN, STaR, or ToT.
@@ 3.3 Theoretical Bounds for GuidedSampling
  • $\mathcal{I}(y;c\;|\;x)$: sample-wise mutual information between $y$ and $c$ conditional on $x$. This represents the additional information contributed by the concept $c$ in predicting $y$.
By intuition, solving a question becomes easier if we know a good problem‑appropriate “hint” for a question. To elaborate on the performance bounds of GuidedSampling, we make the following assumption:
@@
For any “relevant” concept $c\in\mathcal{C}_{r}$, conditioning on it strictly increases the probability of generating a correct solution $y^{*}$. That is, there exists an amplification factor $k_{c}>1$ such that:
$\pi_{base}(y^{*}\;|\;x,c)\geq k_{c}\cdot\pi_{base}(y^{*}\;|\;x)$ (1)
The above assumption is based on the intuition that $\mathcal{I}(y;c\;|\;x)>0$, i.e., any “relevant” concept strictly increases the probability of generating the correct final response. For “irrelevant” concepts ($c\notin\mathcal{C}_{r}$), the assumption doesn’t hold. We also bridge the intuition to above assumption in Appendix A.1. Following the above assumption, we now state our main theorem:
@@
The condition derived in Theorem 1 provides a formal basis for when GuidedSampling outperforms RS. We detail the proof in Appendix A.2. In practice, this condition is satisfied if one or more of the following pathways hold:
@@ 3.4 Post-Training using GuidedSampling
Synthetic data has become an increasingly effective tool for enhancing the reasoning capabilities of LLMs (Gupta et al., 2023; Mitra et al., 2024; Chaudhary et al., 2023). In particular, inference-time algorithms are valuable for generating such data when the correctness of the final solution can be programmatically verified (Zelikman et al., 2022; Arora and Zanette, 2025; Shao et al., 2024). We demonstrate that GuidedSampling can serve not only as an effective inference-time strategy but also as a powerful synthetic data generation mechanism.
@@ 5 Results and Discussion
Figure 4: GuidedSampling forces exploration during inference-time, resulting in $16.01\%$ average pass@k improvement compared to repeated sampling. We observe an average improvement of $21.8\%$ on MATH, $11.87\%$ on GPQA-Diamond, $11.28\%$ on HumanEval, and $3.08\%$ on OlympiadBench. For GuidedSampling, we choose the optimal value of $K$ (from Fig. 5) that maximizes the performance. First row: Llama-3.2-3B-Instruct, Second row: Qwen2.5-3B-Instruct. Third row: Gemma-3-27b-it
@@
The higher performance of GuidedSampling is due to $K$ concepts being generated. In practice, this value is far lower than the number of samples generated (100 in our case). Moreover, as the compute increases (increasing $k$ for pass@k), we observe that the performance gap between Repeated Sampling and GuidedSampling increases in most cases (Fig. 4), suggesting that GuidedSampling benefits more with increased compute. This leads us to believe that when computational resources are sufficient, a small overhead of sequential calls for generating concepts might be a beneficial tradeoff for better performance.
@@
To measure the diversity of candidate solutions, we use Qwen2.5-32B-Instruct (Yang et al., 2024) to extract the core concept or theorem. We then compute the number of distinct concepts generated. On average, RS produces 3.54, 6.72, 2.66, and 3.25 distinct concepts on MATH, GPQA-Diamond, HumanEval, and OlympiadBench, respectively. GuidedSampling produces 3.66, 7.66, 3.87, and 3.81 distinct concepts, improving the diversity by an average of 17.63%. We also found the diversity gains from GuidedSampling are model-specific. We find that Llama-3.2-3B-Instruct generates $3.7\times$ more unique concepts on average compared to Qwen2.5-3B-Instruct, with this gap ranging from $2.82\times$ on GPQA‑Diamond to $5.12\times$ on HumanEval. This suggests that model architecture and pretraining influence the capacity for generating novel reasoning strategies. We show examples of generated concepts in Appendix C.
Figure 5: Pass@50 performance variation with different exploration (number of concepts $K$) and generation (samples per idea $M$) compute allocations, given a fixed total compute of 100 calls ($M=100/K$). Increasing exploration initially helps, but performance declines when the generation budget per idea becomes too small. At $K=0$, GuidedSampling becomes traditional RS. First row: Llama-3.2-3B-Instruct, Second row: Qwen2.5-3B-Instruct. Third row: Gemma-3-27b-it
@@
Performance of Earlier vs Later Concepts
During the concept generation phase of GuidedSampling, concepts are generated iteratively. To determine the contribution of the $k$-th concept across all questions that produced at least $k$ concepts, we analyzed all models and benchmarks mentioned in §4, which contains a total of 1772 questions. We observe a minor decline from $k=1$ to 55 ($19.8\%\rightarrow 16.2\%$). This observation suggests that earlier concepts suggested by the concept generator are better than later ones. However, for concepts with index $k\geq 6$ (i.e., when more exploration is needed), a higher performance variance due to a sharp decrease in coverage is observed. E.g., only 72 out of 1772 questions reach $k=9$, meaning there are fewer samples with $k\geq 6$ concepts. This results in variations in performance, with higher performance being observed for many such cases (e.g., 52.05% performance for $k=14$, due to just 23 instances). Thus, although the earlier concepts are beneficial, later ones ($k\geq 6$) also contribute to increasing performance, but for a small number of instances that require significant exploration. Hence, for the overall success of GuidedSampling, even the later ones are also important, but the earlier ones play a major role. Individual performance values are provided in Appendix 10.
Dependence of GuidedSampling on Well-defined Concepts
Theorem 1 states that for a “relevant” concept $c$, conditioning on $c$ increases the probability of generating the correct solution. However, in domains such as commonsense reasoning, which involve more imprecise, vague, and uncertain knowledge, defining such concepts is difficult. Hence, the condition stated in Theorem 1, i.e., $P(C_{r}|x)>>0$, might not be satisfied. Applying GuidedSampling to Qwen2.5-3B-Instruct on CommonSenseQA (Talmor et al., 2019), a commonsense benchmark. The model is prompted to generate a general idea that could help solve the question (not a task-specific concept, since those are lacking in the commonsense domain). On such domains, GuidedSampling underperforms against Repeated Sampling by 3.28% (pass@50). Based on this, we believe that GuidedSampling has a better chance of succeeding when concepts can be formulated efficiently. More details in Appendix E.
Final Answer Selection via Majority Voting
To select a final solution after sampling multiple times, we use the majority voting technique, where the most common solution is selected as the final answer. GuidedSampling achieves an average accuracy of 35.87% compared to Repeated Sampling (32.80%) and Tree-of-Thought (26.26%). Detailed accuracies in Appendix E.3.
@@ 6 Conclusions
Limitations & Future Work
While our method is successful in improving the diversity of solutions generated by LLMs, it represents an early step in this area and has some limitations, including but not limited to the following:
Limited model coverage
While our evaluation spans three open-source models and a limited study on GPT-4o-mini and Phi-4-mini-instruct, applying GuidedSampling to proprietary models (e.g., GPT-5, Gemini-2.5-Pro) remains unexplored due to high inference costs and lack of training access. Extending the method to state-of-the-art models is an important direction for future work.
Exploration cost vs. effectiveness trade-off
Although our method improves diversity, the optimal balance between the number of concepts ($K$) and samples per concept ($M$) under a fixed compute budget remains task-specific. Developing adaptive strategies for this trade-off is a promising area.
Generality across domains
Our work demonstrates promising results in mathematical, scientific, and code generation domains. However, further evaluation is needed to understand how well GuidedSampling generalizes to more diverse domains such as legal reasoning, medical, or discovery.
Concept generation quality
The success of GuidedSampling depends on the quality and diversity of the generated concepts. Although some “irrelevant” concepts can help (Appendix G), investigating techniques to improve/verify the relevance of these concepts (e.g., through external tools or feedback mechanisms) can enhance overall effectiveness.
Better Verifier
GuidedSampling has an exploration phase, which forces the model to explore multiple concepts, increasing diversity. This can lead to multiple final solutions. While this increases pass@k, building a robust verifier that can select a final solution, even if it is in the minority, remains a challenging future task.
Reprodubility Statement
To ensure the reproducibility of our results, we release the source code and data through our GitHub repo https://github.com/DivijH/sampling_inference. We also note that LLMs are inherently probabilistic in nature, and some results may vary upon each run. We hope our code and data aid in future research.
Ethics Statement
In accordance with ICLR policy, we disclose that AI assistants, specifically Grammarly for grammar correction and ChatGPT for sentence restructuring and paraphrasing, were utilized during the preparation of this manuscript. The authors have reviewed, edited, and take full responsibility for all final content presented in this paper.
Acknowledgement
We thank the anonymous reviewers for their constructive suggestions. We extend our gratitude to the Research Computing (RC) and Enterprise Technology at ASU for providing computing resources for experiments.
Appendix A Theoretical Proofs
A.1 From Intuition to Assumption 1
Assumption 1 stems from the intuition that any “relevant” concept helps in answering a given question, i.e., the concept adds more information which is useful. This is represented as sample-wise information between $y$ and $c$ conditioned on $x$.
$\mathcal{I}(y;c\;|\;x)>0$
$\log\pi_{base}(y^{*}\;|\;x,c)\;-\;\log\pi_{base}(y^{*}\;|\;x)>0$
$\log\pi_{base}(y^{*}\;|\;x,c)>\log\pi_{base}(y^{*}\;|\;x)$
$\pi_{base}(y^{*}\;|\;x,c)\geq k_{c}\cdot\pi_{base}(y^{*}\;|\;x)$ (3)
This is the stated assumption 1.
A.2 Proof of Theorem 1
@@ B.2 Concept Extraction Prompt
Concept Extraction Prompt You are ConceptTagger, an expert that maps a worked-out solution (chain-of-thought or final answer) to the most specific mathematical or logical concept that makes the solution possible.
@@
B.3 CAA Prompt
CAA Data I have a few ideas to solve this problem.
@@ Appendix E More results using GuidedSampling
E.1 Results for more LLMs
In this section, we showcase some results on additional models. As mentioned in §4, we generate 100 candidate solutions for each instance. We provide results on Phi-4-mini-instruct (Abouelenin et al., 2025), GPT-4o-mini (Hurst et al., 2024), and Gemma-3-27b-it (Team et al., 2025). Due to limited resource constraints, we limit the proprietary model to just the MATH (Hendrycks et al., 2021) benchmark. Table 6 and 7 show the pass@50 results for these models along with the observed diversity as extracted by Qwen-3.2-32B-Instruct (Yang et al., 2024). Diversity is measured by the average number of concepts for each instance.
@@ E.1 Results for more LLMs
Table 7: pass@50 performance of Gemma-3-27b-it
Benchmark Repeated Sampling GuidedSampling
MATH 81.00% 83.80%
GPQA-Diamond 70.20% 92.42%
HumanEval 83.54% 94.51%
OlympiadBench 33.19% 43.52%
E.2 Results on CommonSenseQA
Results for Qwen2.5-3B-Instruct on CommonSenseQA are reported in Table 8. The prompts used don’t specify a task-specific definition of concepts. Prompts are as follows:
CommonSenseQA Initial Concept Generation You are a helpful assistant. Your task is to state a concept that is relevant and useful for answering the question.
QUESTION:
{ele[‘question’]}
Provide the concept that is most directly applicable to answering the question. Do not answer the original question.
CommonSenseQA Subsequent Concept Generation You are a helpful assistant. You will be presented with a question and a list of concepts that have already been proposed as potentially useful for answering the question. Your task is to provide a *new* and *different* concept that is relevant and useful for answering the question.
QUESTION:
{ele[‘question’]}
EXISTING CONCEPTS:
{ideas_text}
Provide the concept that is most directly applicable to answering the question. Do not answer the original question. If no new, distinct, and useful concept can be identified, respond with “No additional concepts found.”
Table 8: pass@50 performance of Qwen2.5-3B-Instruct on CommonSenseQA. RS: Repeated Sampling, GS: GuidedSampling
Repeated Sampling GuidedSampling
98.94% 95.66%
E.3 Majority Voting Results
Table 9 shows the overall accuracies of Majority Voting applied on top of Repeated Sampling, GuidedSampling, and Tree-of-thought. Out of the 8 different settings, GuidedSampling achieves better accuracy in 4 of them, and a higher average performance as well.
Table 9: Accuracy of models on benchmarks using majority voting.
Benchmark Model Repeated Sampling GuidedSampling Tree-of-thought
MATH Llama-3.2-3B-Instruct 50.40% 43.40% 45.80%
GPQA-Diamond Llama-3.2-3B-Instruct 23.23% 23.23% 19.19%
HumanEval Llama-3.2-3B-Instruct 20.12% 45.12% 25.61%
OlympiadBench Llama-3.2-3B-Instruct 17.47% 18.35% 12.75%
MATH Qwen2.5-3B-Instruct 51.20% 64.20% 45.40%
GPQA-Diamond Qwen2.5-3B-Instruct 20.71% 20.20% 7.07%
HumanEval Qwen2.5-3B-Instruct 56.71% 50.61% 39.02%
OlympiadBench Qwen2.5-3B-Instruct 22.53% 21.87% 15.27%
Average - 32.80% 35.87% 26.26%
@@ Appendix F Finetuning Setup
To determine whether the model trained using CAA trajectories experiences any collapse, we use one common observation: a collapsed model can repeat tokens indefinitely without generating an end-of-sequence token during inference. While model collapse has been studied to occur for several reasons (Shumailov et al., 2024; Gerstgrasser et al., 2024), checking for repeated tokens can indicate whether collapse happens or not.
To validate this, we run the base model and the model trained on CAA trajectories on HumanEval with 10 candidate solutions and check the “finish_reason”  [^2] https://docs.vllm.ai/en/latest/api/vllm/outputs/#vllm.outputs.CompletionOutput.finish_reason after generation. Both the base model and the model trained using CAA trajectories return with the finish reason of “stop”, indicating that the model produced the end-of-sequence token.
Moreover, our evaluation, as described in §4, requires the model to generate responses in a particular format (refer to Appendix B.3 for the prompt). Since we did see improvements in pass@k values, we can be assured that the model followed the format closely.
Based on these two insights, we can say with some evidence that no model collapse occurred while training with CAA trajectories.
@@ Appendix H Algorithm for GuidedSampling
Appendix I Performance variation for $k$-th Concept
Figure 6: Pass@50 performance variation for $k$-th concept averaged across all benchmarks and mdoels mentioned in §4.
Table 10: pass@50 performance and the number of instances for the $k$-th concept generated in GuidedSampling across all benchmarks and models, resulting in a total of 1772 instances.
Concept Index Avg. Correct Solutions (in %) Number of Instances
1 19.76 % 1772 (100.00%)
2 19.93 % 1646 (92.89%)
3 18.54 % 1473 (83.13%)
4 17.57 % 1193 (67.33%)
5 16.19 % 819 (46.22%)
6 18.71 % 178 (10.05%)
7 17.98 % 126 (7.11%)
8 15.13 % 89 (5.02%)
9 22.25 % 72 (4.06%)
10 19.98 % 59 (3.33%)
11 18.34 % 47 (2.65%)
12 26.31 % 39 (2.20%)
13 18.96 % 28 (1.58%)
14 52.05 % 23 (1.30%)
15 43.12 % 16 (0.90%)
16 42.14 % 14 (0.79%)
17 35.50 % 8 (0.45%)
18 30.86 % 7 (0.40%)
19 44.80 % 5 (0.28%)
20 0.00 % 1 (0.06%)
Here we detail the individual performance of the $k$-th concept across every model and benchmark. Fig. 6 illustrates the performance for every concept. As discussed in §5. Since later concepts have fewer instances, we see a huge variation in performance. Table 10 shows the detailed performance and number of instances for all concepts.
Appendix J Latency of Inference-Time Algorithms
Figure 7: Pass@50 performance against the number of LLM calls for different inference-time algorithms averaged across all models and benchmarks.
Figure 7 shows the relationship between the number of LLM calls and pass@50 performance for Repeated Sampling (RS), GuidedSampling (GS), and Tree-of-Thought (ToT). All results are averaged across all models and benchmarks. We found that GuidedSampling (pass@50=60.2 with 104.75 calls) outperforms both Repeated Sampling (pass@50=48.2 with 100 calls) and Tree-of-Thought (pass@50=37.1 with 154 calls), while being more efficient than Tree-of-Thought.